# A New Look at Fiducial Inference

### v1

@@ 4.2 An exponential example
The Fisher fiducial density function is given by an inverse gamma density with parameters $(n,n\theta_{n})$. This is also the posterior arising from the objective prior proportional to $1/\theta$. For the details, the statistic $t_{n}=\sum_{i=1}^{n}x_{i}$ is a realization of a gamma random variable with parameters $(n,1/\theta)$ so the fiducial density is
[…]
which becomes the said inverse gamma density.
Our fiducial posterior is based on sampling $x_{n+1:\infty}$ via $x_{m+1}$ is taken from $f(x\mid t_{m})$ and $t_{m}=\bar{x}_{m}$. Hence, a sample from the posterior is
$t_{\infty}=t_{n}\,\prod_{i=n}^{\infty}\frac{i+z_{i}}{i+1},$
where the $(z_{i})$ are realizations of independent standard exponential random variables. So $t_{\infty}=t_{n}\,z$ where
$z=\prod_{i=n}^{\infty}\frac{i+z_{i}}{i+1},$
which is a realization of a finite r.v.
@@
It can be shown by standard arguments that $\sum_{m}(1-a_{m})<\infty$. The term of order $1/m$ disappears since $\frac{1}{2}\int_{0}^{\infty}(1-z)e^{-z}\rm{d}z=0$. Hence, the product martingale defined by
$M_{n}=\prod_{i=1}^{n}\frac{i+Z_{i}}{i+1},$
where $Z_{1},Z_{2},\dotsc$ are independent standard exponential random variables, converges with probability one to $M_{\infty}$ and ${\rm E}(M_{\infty})=1$ and $M_{n}\to M_{\infty}\,\,\mbox{in }L_{1}.$ In the next section we will look at more general theory for convergence.
@@ 4.3 Weibull model
With a sample of size $n$ one will have $x_{i}=z_{i}^{1/\theta}$, $i=1,\ldots,n$ where the same rules apply to each $(x_{i},z_{i})$ as before. The problem now is that due to a lack of a statistic corresponding to the parameter, a least squares solution is provided as part of the inversion procedure; i.e. $\theta$ minimizes
$\sum_{i=1}^{n}\left(\log x_{i}-\theta^{-1}\log u_{i}\right)^{2}$
so
@@
Assuming we have a sample of size $n=50$ with $\theta_{n}=3$, Fig. 4 shows a sample of $\theta$s from our fiducial distribution, by running each sequence up to a size $m=1000$.
@@ 4.4 Uniform model
Consider the uniform distribution on $[0,\theta]$. In this setting, a suitable estimate which is unbiased and a sufficient statistic is $t_{n}=(n+1)/n\,x_{(n)}$, where $x_{(n)}=\max\{x_{1},\dotsc,x_{n}\}$. According to the sampling scheme where $x_{m+1}=t_{m}\,u_{m}$ where $u_{m}$ is a realization of a uniform random variable $U_{m}$ on $[0,1]$ independent of $T_{m}$, we have that
@@
[…]
@@
so that ${\rm E}(X_{m})=1,$, and
$T_{m+1}=T_{n}\prod_{j=n}^{m}(\mu_{j}X_{j}),$
for every $m\geq n$. Moreover,
@@
As a consequence, $T_{m}$ converges almost surely to a positive random variable $T_{\infty}$ as $m$ diverges to infinity such that
${\rm E}(T_{\infty}\mid T_{n}=t_{n})=t_{n}\,\prod_{m=n}^{\infty}\mu_{m}.$
This means that the expectation of our fiducial distribution is between $x_{(n)}$ and $x_{(n)}\,e^{1/(2n)}$. Instead, the mean of Fisher fiducial distribution is $n/(n-1)x_{(n)}$. Of course, the difference between these two estimators is negligible for large $n$.
Our approach allows us to find the fiducial distribution of both parameters of the uniform distribution. Consider the uniform distribution on $[a-b,a+b]$. In this setting, unbiased estimates for the parameters $a$ and $b$ based on sufficient statistics are
@@
and therefore also the random sequence $(A_{m})$, whose realization $(a_{m})$ are generated recursively by (16), converges almost surely in virtue of Theorem 1.
@@ 5 Discussion
Our approach to fiducial inference is general and based on the idea that we can generate a sequence of statistics, say $(T_{m})_{m>n}$, where $T_{n}=t_{n}$ is observed. It is desirable to have
$T_{m+1}=H_{m}(T_{m},z_{m})$
for some sequence of functions $(H_{m})$, where the $(Z_{m})$ are a sequence of independent variables and independent of the $(T_{m})$. There are many cases when this holds. For example, if $X_{m+1}=G(T_{m},Z_{m})$ and $T_{m+1}=G_{m}(T_{m},X_{m+1})$.
@@
where the $(Z_{m})_{m>n}$ are independent standard normal random variables. Hence, in this case, we can write
$F_{m+1}=H_{m}(F_{m},Z_{m}),\quad m\geq n.$
[…]

### v2

@@ 4.2 An exponential example
The Fisher fiducial density function is given by an inverse gamma density with parameters $(n,n\theta_{n})$. This is also the posterior arising from the objective prior proportional to $1/\theta$. For the details, the statistic $t_{n}=\sum_{i=1}^{n}x_{i}$ is a realization of a gamma random variable with parameters $(n,1/\theta)$ so the fiducial density is $\pi(\theta\mid x_{1:n})=-\frac{\partial}{\partial\theta}\int_{0}^{t_{n}/\theta}s^{n-1}e^{-s}\,ds/\Gamma(n)$ which becomes the said inverse gamma density.
Our fiducial posterior is based on sampling $x_{n+1:\infty}$ via $x_{m+1}$ is taken from $f(x\mid t_{m})$ and $t_{m}=\bar{x}_{m}$. Hence, a sample from the posterior is given by $t_{\infty}=t_{n}\,\prod_{i=n}^{\infty}(i+z_{i})/(i+1),$ where the $(z_{i})$ are realizations of independent standard exponential random variables. So $t_{\infty}=t_{n}\,z$ where $z=\prod_{i=n}^{\infty}(i+z_{i})/(i+1),$ which is a realization of a finite r.v.
@@
It can be shown by standard arguments that $\sum_{m}(1-a_{m})<\infty$. The term of order $1/m$ disappears since $\frac{1}{2}\int_{0}^{\infty}(1-z)e^{-z}\rm{d}z=0$. Hence, the product martingale defined by $M_{n}=\prod_{i=1}^{n}(i+Z_{i})/(i+1),$ where $Z_{1},Z_{2},\dotsc$ are independent standard exponential random variables, converges with probability one to $M_{\infty}$ and ${\rm E}(M_{\infty})=1$ and $M_{n}\to M_{\infty}\,\,\mbox{in }L_{1}.$ In the next section we will look at more general theory for convergence.
@@ 4.3 Weibull model
With a sample of size $n$ one will have $x_{i}=z_{i}^{1/\theta}$, $i=1,\ldots,n$ where the same rules apply to each $(x_{i},z_{i})$ as before. The problem now is that due to a lack of a statistic corresponding to the parameter, a least squares solution is provided as part of the inversion procedure; i.e. $\theta$ minimizes the least squares $\sum_{i=1}^{n}\left(\log x_{i}-\theta^{-1}\log u_{i}\right)^{2}$ so
@@
[…]
@@ 4.4 Uniform model
Consider the uniform distribution on $0,\theta]$. In this setting, a suitable estimate which is unbiased and a sufficient statistic is $t_{n}=(n+1)/n\,x_{(n)}$, where $x_{(n)}=\max\{x_{1},\dotsc,x_{n}\}$. According to the sampling scheme where $x_{m+1}=t_{m}\,u_{m}$ where $u_{m}$ is a realization of a uniform random variable $U_{m}$ on $[0,1]$ independent of $T_{m}$, we have that
@@
Fig. 5 shows our fiducial distribution as a histogram, taking $n=50$ and the $x_{(n)}=0.947$, with the true $\theta$ being 1. The Fisher fiducial distribution is shown as a line and is the Pareto density $\pi(\theta\mid x_{(n)})=nx_{(n)}^{n}/\theta^{n+1}\,1(\theta>x_{(n)})$.
@@
so that ${\rm E}(X_{m})=1,$, and $T_{m+1}=T_{n}\prod_{j=n}^{m}(\mu_{j}X_{j}),$ for every $m\geq n$. Moreover,
@@
As a consequence, $T_{m}$ converges almost surely to a positive random variable $T_{\infty}$ as $m$ diverges to infinity such that ${\rm E}(T_{\infty}\mid T_{n}=t_{n})=t_{n}\,\prod_{m=n}^{\infty}\mu_{m}.$ This means that the expectation of our fiducial distribution is between $x_{(n)}$ and $x_{(n)}\,e^{1/(2n)}$. Instead, the mean of Fisher fiducial distribution is $n/(n-1)x_{(n)}$. Of course, the difference between these two estimators is negligible for large $n$.
Our approach allows us to find the fiducial distribution of both parameters of the uniform distribution. Consider the uniform distribution on [ab,a+b][a-b,a+b]. In this setting, unbiased estimates for the parameters $a$ and $b$ based on sufficient statistics are
@@
and therefore the random sequence $(A_{m})$, whose realization $(a_{m})$ are generated recursively by (17), converges almost surely by virtue of Theorem 1.
5 Nonlinear regression model
Consider the following regression model, $Y=g(X,Z,\theta)$ where $g$ is a known regression function, $Y\in\mathbb{R}$ is the response, $X\in\mathbb{R}^{p}$ are the covariates, $\theta\in\mathbb{R}^{d}$ is the unknown parameter, $Z\in\mathbb{R}$ is the error following a known distribution and independent of $X$.
Given observed data $(x_{1},y_{1}),\dotsc,(x_{n},y_{n})$, an estimate $\hat{\theta}_{n}$ of $\theta$ is obtained, for example by minimizing $\sum_{i=1}^{n}L(\theta,x_{i},y_{i})/n$, where
$L(\theta,x,y)=(y-\mu(x,\theta))^{2}$ (19)
and $\mu(x,\theta)={\rm E}(g(x,Z,\theta)).$ In other words,
$\hat{\theta}_{n}=\arg\min_{\theta\in\mathbb{R}^{d}}\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\mu(x_{i},\theta))^{2}$
We model the sequence of future observations as an asymptotically exchangeable random sequence $(X_{n+1},Y_{n+1}),(X_{n+2},Y_{n+2}),\dotsc$ with the aim to collect posterior samples of the random vector $T$ such that:
$T=\arg\min_{\theta\in\mathbb{R}^{d}}\int_{\mathbb{R}^{p}\times\mathbb{R}}(y-\mu(x,\theta))^{2}\,\nu_{\infty}({\rm d}x,{\rm d}y)$ (20)
almost surely (provided that $T$ exists and is almost surely unique), where $\nu_{\infty}$ is the almost sure weak limit of the empirical measure $\nu_{m}$, as $m$ diverges to infinity, $\nu_{m}=m^{-1}\sum_{i=1}^{m}\delta_{(X_{i},Y_{i})}.$ From (19), for $j=1,\dotsc,d$,
$L^{\prime}_{\theta_{j}}(\theta,x,y)=2(y-\mu(x,\theta))\mu^{\prime}_{\theta_{j}}(x,\theta),$
future observations are obtained through the following recursive update based on stochastic gradient descent:
$Y_{m+1}$ $=g(X_{m+1},Z_{m+1},T_{m})$ (21)
$T_{m+1,j}$ $=T_{m,j}-\frac{(Y_{m+1}-\mu(X_{m+1},T_{m}))\mu^{\prime}_{\theta_{j}}(X_{m+1},T_{m})}{(m+1)\varphi(X_{m+1},T_{m})}$ (22)
for $m=n,n+1,\dotsc$, where:
$\varphi(x,\theta)=\sqrt{\mu_{2}(x,\theta)-\mu(x,\theta)^{2}}\max_{j=1,\dotsc,d}\lvert\mu^{\prime}_{\theta_{j}}(x,\theta)\rvert$,
$\mu_{2}(x,\theta)={\rm E}(g^{2}(x,Z,\theta)),$ $Z_{n+1},Z_{n+2},\dotsc$ are i.i.d. and distributed as $Z$, and $X_{n+1},X_{n+2},\dotsc$ are obtained through the Bayesian bootstrap (starting from $x_{1},\dotsc,x_{n}$), and independent of $Z_{n+1},Z_{n+2},\dotsc$.
Theorem 4.
Assume that $\mu(x,\theta)$ is invertible as a function of $\theta$, for every $x\in\mathbb{R}^{p}$. If (21) and (22) hold then $\lim_{m\to\infty}T_{m}=T$, almost surely, where $T$ satisfies (20). Moreover, the random sequence $(X_{n+1},Y_{n+1}),(X_{n+2},Y_{n+2}),\dotsc$ is asymptotically exchangeable.
Proof.
For each $j=1,\dotsc,d$, $(T_{n+1,j},T_{n+2,j},\dotsc)$ is a martingale with respect to the filtration $(\mathscr{G}_{n+1},\mathscr{G}_{n+2},\dotsc)$ where
$\mathscr{G}_{m}=\sigma(Z_{n+1},\dotsc,Z_{m},Y_{n+1},\dotsc,Y_{m},X_{n+1},\dotsc,X_{m+1})$
for $m=n+1,n+2,\dotsc$ Moreover, by (22), we have that:
$\begin{split}\textrm{Var}(T_{m+1,j})&=\textrm{Var}({\rm E}(T_{m+1,j}\mid\mathscr{G}_{m}))+{\rm E}(\textrm{Var}(T_{m+1,j}\mid\mathscr{G}_{m}))\\ &\leq\textrm{Var}(T_{m,j})+\frac{1}{(m+1)^{2}}\end{split}$
so that for every $m=n+1,n+2,\dotsc$, and $j=1,\dotsc,d$,:
$\textrm{Var}(T_{m,j})\leq\sum_{\ell=n+1}^{m}\frac{1}{\ell^{2}}<\sum_{\ell=n+1}^{\infty}\frac{1}{\ell^{2}}<\infty$
and therefore there exists a random variable $T_{j}$ such that $\lim_{m\to\infty}T_{m,j}=T_{j}$, almost surely, for $j=1,\dotsc,d$.
By Lemma 8.2 in Aldous (1985), to prove asymptotic exchangeability we need to prove that there exists a random probability measure $\nu$ on $\mathbb{R}^{p}\times\mathbb{R}$ such that for every bounded and continuous function $h_{1}:\mathbb{R}^{p}\to\mathbb{R}$ and $h_{2}:\mathbb{R}\to\mathbb{R}$ such that:
$\lim_{m\to\infty}{\rm E}(h_{1}(X_{m+1})\,h_{2}(Y_{m+1})\mid X_{1:m},Y_{1:m})=\int_{\mathbb{R}^{p}\times\mathbb{R}}h_{1}(x)\,h_{2}(y)\,\nu({\rm d}x,\,{\rm d}y)$ (23)
To this aim, note that:
$\begin{split}{\rm E}(h_{1}(X_{m+1})\,&h_{2}(Y_{m+1})\mid X_{1:m},Y_{1:m})\\ &={\rm E}(h_{1}(X_{m+1})\,\zeta(X_{m+1},T_{m})\mid X_{1:m},Y_{1:m})\end{split}$ (24)
where $\zeta(x,\theta)={\rm E}(h_{2}\circ g(x,Z,\theta))$. At this stage, recall that the sequence $X_{n+1},X_{n+2},\dotsc$ obtained through the Bayesian bootstrap is exchangeable. Let $X_{0}$ be a random variable independent of $Z_{1:\infty}$ and of $Y_{1:\infty}$ such that the sequence $(X_{0},X_{n+1},X_{n+2},\dotsc)$ is exchangeable. In this way, the conditional distribution of $X_{0}$ given $X_{1:m}$ is the same of $X_{m+1}$ given $X_{1:m}$. Therefore, (24) becomes:
${\rm E}(h_{1}(X_{m+1})\,h_{2}(Y_{m+1})\mid X_{1:m},Y_{1:m})={\rm E}(h_{1}(X_{0})\,\zeta(X_{0},T_{m})\mid X_{1:m},Y_{1:m})$ (25)
At this stage, recall that $\lim_{m\to\infty}T_{m}=T$, almost surely. By the dominated convergence theorem for conditional expectations (see for instance Theorem 4.6.10 in Durrett, 2019), (25) yields:
$\lim_{m\to\infty}{\rm E}(h_{1}(X_{m+1})\,h_{2}(Y_{m+1})\mid X_{1:m},Y_{1:m})={\rm E}(h_{1}(X_{0})\,\zeta(X_{0},T)\mid X_{1:\infty},Y_{1:\infty})$
[…]
To conclude the proof we need to show that $T$ satisfies (20), namely:
$T=\arg\min_{\theta\in\mathbb{R}^{d}}{\rm E}(\tilde{Y}-\mu(\tilde{X},\theta))^{2}.$ (26)
If the random vector $\tilde{T}$ is a solution of (26), we have that
$\begin{split}{\rm E}(\tilde{Y}-\mu(\tilde{X},\tilde{T}))^{2}&={\rm E}({\rm E}(\tilde{Y}-\mu(\tilde{X},\tilde{T})^{2}\mid\tilde{X}))\\ &={\rm E}(\textrm{Var}(\tilde{Y}\mid\tilde{X}))+{\rm E}(\mu(\tilde{X},T)-\mu(\tilde{X},\tilde{T}))^{2}\end{split}$
[…]
5.1 Logistic regression
To illustrate the above family of models, we consider a logistic regression model, i.e.
$P(Y=1\mid x)=\frac{e^{x^{\prime}\beta}}{1+e^{x^{\prime}\beta}}$
where $x$ is a $p$ dimensional vector of covariates and $\beta$ a $p$ dimensional vector of parameters. The function $g(X,Z,T)$ can be represented as $g(x,Z,\beta)=1(Z<P(Y=1\mid x))$ where $Z$ is an independent uniform random variable on $(0,1)$ and $\mu(x,\beta)=P(Y=1\mid x)$. It is easy to check that
$\mu_{j}^{\prime}(x,\beta)=\frac{x_{j}\,e^{x^{\prime}\beta}}{(1+e^{x^{\prime}\beta})^{2}}$
and
$\phi(x,\beta)=\sqrt{\mu(x,\beta)(1-\mu(x,\beta))}\,\frac{e^{x^{\prime}\beta}}{(1+e^{x^{\prime}\beta})^{2}}\,\max_{j}|x_{j}|.$
Hence, one we have an estimator based on an observed sample of size $n$, say $\beta_{n}$, we can implement the algorithm generated by (21) and (22) to generate samples from the fiducial posterior distribution.
Figure 6: Fiducial posterior for regression parameters for the Titanic dataset
We illustrate with the Titanic data set, which is available on CRAN, for example. The data consists of $n=887$ individuals and records their survival or otherwise along with covariate information including age, sex (male=1), cost of ticket, and class of travel, of which there are three types, 1,2 and 3. The covariates are all non-negative and standardized to lie between 0 and 1. We also include an intercept term, and so there are 5 unknown parameter values, which are estimated as $\widehat{\beta}=(4.98,-2.74,-4.79,2.28,-3.22).$ We then ran the algorithm generated by (21) and (22) for a length of size $n+1000$ with the starting point being $n+1$ and starting with $\beta_{n}=\widehat{\beta}$. A repetition of this over a 1000 runs leads to fiducial posterior samples from which we can construct histograms; see Fig. 6.
@@ 6 Discussion
Our approach to fiducial inference is general and based on the idea that we can generate a sequence of statistics, say $(T_{m})_{m>n}$, where $T_{n}=t_{n}$ is observed. It is desirable to have $T_{m+1}=H_{m}(T_{m},z_{m})$ for some sequence of functions $(H_{m})$, where the $(Z_{m})$ are a sequence of independent variables and independent of the $(T_{m})$. There are many cases when this holds. For example, if $X_{m+1}=G(T_{m},Z_{m})$ and $T_{m+1}=G_{m}(T_{m},X_{m+1})$.
@@
where the $(Z_{m})_{m>n}$ are independent standard normal random variables. Hence, in this case, we can write $F_{m+1}=H_{m}(F_{m},Z_{m}),\quad m\geq n.$ The limit $F_{\infty}$ is a sample from the Doob fiducial distribution. To adjust to a statistic $T_{m}$ it would be typical that $T_{m}=\int g(x)\,dF_{m}(x)$ in which case, grouping $S_{m}=(T_{m},F_{m})$, we have $S_{m+1}=H_{m}(S_{m},z_{m}).$